# references.bib

### v3

@@ 3 Area estimation
3.1 Simulation: Maize area in Iowa
@@
In each trial, we sample $n=100$ “ground truth” points uniformly at random from the $N=100000$ map product points. Initially, we set the map product and ground truth equal to each other at the 100 ground truth locations. We then experiment with degrading the map product quality as follows. We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), classical, and PPI estimators. We average results over 100 trials.
@@
We define adding unbiased noise at level $p$ to the map product as resampling each of the 100000 map product points with probability $p$ from the Bernoulli(0.35) distribution. In other words, about $p$ fraction of the points are randomly labeled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 and non-maize with probability 0.65. We experiment with noise probabilities $p$ from 0 to 1 at increments of 0.1.
@@ Results
As we add increasing amounts of random noise to the map, all confidence intervals for maize area remain centered at the true maize fraction of 0.35 (Figure 5). This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same. The classical confidence interval stays constant because it only uses ground truth points. The map-only confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product. However, the map-only confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
When the noise probability is low, the PPI confidence intervals are significantly smaller than the classical intervals. This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error. As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the classical confidence interval at noise probability $p=1$. (At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth samples.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$). Thus, the effectiveness of PPI is dependent on the quality of the map product. The post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI.
@@
The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 6. Again, the classical confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points. As bias increases, the fraction of map product points labeled as “maize” increases. As a result, the map-only estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35. The map-only confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product. In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product. The PPI confidence intervals are smaller than the classical intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level). Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points. Again, the post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
We also plot the maize fraction confidence intervals against model accuracy for both the noise and bias experiments (Figure 7). As noise increases, model accuracy decreases, so low model accuracy is associated with wide PPI and post-stratified confidence intervals. Thus, the left panel of Figure 7 resembles an inverted version of the left panel of Figure 5.
Figure 7: Simulation estimating maize fraction 95% confidence intervals at different map product accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points. Post-stratified and PPI confidence intervals are overlapping.
As bias increases and the fraction of map product points labeled as “maize” increases, model accuracy decreases linearly. (To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as model accuracy increases. In the right panel of Figure 7, a model accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35. This shows that even a map with high model accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
3.2 Deforestation area in Brazilian Amazon
@@
We create a binary deforestation map product using all $N=4.4\times 10^{9}$ pixels from the NASA Global Forest Cover Change (GFCC) map [townshend2016global] (at 30 m resolution) in the study area (Figure 8). The GFCC product is derived from Landsat 5 and Landsat 7 surface reflectance imagery, and includes maps for tree cover percentage in each pixel in five-year increments from 2000 to 2015. We consider a map pixel deforested if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$ (Appendix B).
We use a ground truth Amazon deforestation and forest disturbance dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth between 1995 and 2017. The dataset records the years corresponding to deforestation, degradation, and natural disturbance events at each ground truth point. We use this to construct a binary ground truth dataset restricted to Brazil, in which we consider a point deforested if it experienced a deforestation event between 2000 and 2015. The original dataset is obtained by stratified random sampling, using a stratification map (also created by [bullock2020satellite]) that contains 7 strata related to deforestation status. To generate a classical estimate baseline, we construct a “uniformly random sampled” ground truth dataset by downsampling the points in each of the strata to match its area proportion. For a fair comparison, we also downsample the original stratified ground truth dataset to match the size of the “uniformly random” dataset. This results in two ground truth datasets of $n=1386$ points each, corresponding to uniform random sampling and stratified random sampling. Both ground truth datasets are used with PPI to obtain confidence intervals.
@@ Results
Figure 9: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points. The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than classical estimation, with Stratified PPI having the best performance.
Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015. The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than classical estimation (Figure 9). The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth samples. The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the classical method by a factor of about $2.4$. (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth samples.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth. In particular, the stratified estimator is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
@@
4 Estimating regression coefficients
In this section, we apply our uncertainty quantification methods to estimating regression coefficients for associative relationships between a remotely sensed outcome and multiple covariates. Regressions are commonly used in the environmental sciences and economics to identify relationships — associative or causal — between covariates and a binary outcome (logistic regression) or between covariates and a continuous outcome (linear regression). Logistic regression coefficients give the change in the log odds ratio of a binary outcome corresponding to a one-unit increase in a covariate. Linear regression coefficients give the change in the value of a continuous outcome corresponding to a one-unit increase in a covariate. Thus, we can interpret these coefficients as effect sizes of the covariates.
In particular, we estimate (1) logistic regression coefficients for Brazilian Amazon deforestation in 2000–2015, and (2) linear regression coefficients for US tree cover in 2021. We compare the classical method, map-only estimator, PPI, stratified estimator, and stratified PPI. Our methods only account for uncertainty in the remotely sensed target variable, while ignoring uncertainty in the covariates; we leave addressing the latter to future work.
4.1 Deforestation logistic regression
Problem Formulation
We estimate the logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates: distance from major roads, distance from major rivers, elevation, and slope (Figure 10). We use logistic regression because the outcome variable (deforested or non-deforested) is binary.
Given $n$ data points, the logistic regression model is
$\log\frac{p}{1-p}=\theta_{\text{road}}\textit{x}_{\text{road}}+\theta_{\text{river}}\textit{x}_{\text{river}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
where $p$ is a $n\times 1$ vector of probabilities, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar logistic regression coefficient for covariate $j$. We wish to estimate the $\theta_{j}$. A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the log-odds of the outcome variable.
We use the same Brazilian Amazon deforestation datasets as the previous section, with $n=1386$ ground truth points and $N=963548$ uniformly randomly sampled map product points. For the covariates, we use a map of Brazilian federal roads in 2000 [de2021transport] and the WWF HydroSHEDS Free Flowing Rivers Network map [grill2019mapping] (we include all rivers of order at most 3) to compute the distance from each data point to the nearest major road and major river. We obtain elevation and slope from the NASA Digital Elevation Model [jpl2020] at 30 m resolution. Prior work has shown that deforestation in the Brazilian Amazon is more likely to occur near major roads and navigable rivers due to human settlement and economic activities along these transportation networks [laurance2002predictors, kirby2006future, rosa2013predictive]. Elevation and slope have also been associated with forest cover change; for example, deforestation is generally less likely on sloped terrain [sandel2013human]. We do not include a comprehensive list of covariates that could influence deforestation, because the primary purpose of this regression is to illustrate uncertainty quantification rather than explain deforestation.
Figure 10: Brazilian Amazon deforestation logistic regression datasets. We regress the deforestation binary target variable (left) against four continuous covariates (right): distance from roads, distance from rivers, elevation, and slope.
Results
The 95% confidence intervals for the four logistic regression coefficients are shown in Figure 11.
The general degree of uncertainty in estimating the coefficients from our non-map-only methods are as follows. The log odds of deforestation change by about .$-.016$ to .$-.010$ for each one-kilometer increase in distance from roads, .$-.008$ to .$-.001$ for each one-kilometer increase in distance from rivers, $-3$ to $0.8$ for each one-kilometer increase in elevation, and $-0.28$ to $-0.11$ for each one-degree increase in slope. Thus, deforestation is negatively associated with all covariates except for elevation. The elevation coefficient confidence intervals intersect with zero for all non-map-only methods except for stratified PPI, suggesting that there might be no association between elevation and deforestation after accounting for the other three variables.
The PPI methods result in the smallest confidence intervals, with effective sample size outperforming the classical method by a factor of 1.1 to 2.5 depending on the covariate. The most significant improvement is for the slope covariate. Unlike in our previous area estimation examples, the stratified estimator (i.e., weighted logistic regression with weights determined by the strata) is not effective and even reduces effective sample size relative to the classical method when estimating the coefficients for elevation and distance from rivers. Similarly, stratified PPI has slightly worse performance than PPI for all covariates except distance from roads. The strata were designed to minimize variance in deforestation status, which does not necessarily correspond to better estimates of regression coefficients for the covariates.
For slope and distance from rivers, the map-only estimated coefficient computed from the map product falls outside the confidence intervals for all methods that use ground truth points. The map-only coefficient for slope is about 0, and the map-only coefficient for distance from rivers is about $-0.010$. In particular, for the slope coefficient, the methods that use ground truth points find an association that is statistically significantly different from zero. However, if we had used only the remote sensing map, we would not have discovered the negative relationship between slope and deforestation. Thus, using only the map product could result in misleading conclusions due to biases in the deforestation map product that are correlated with covariates (e.g., the map product overestimates deforestation at locations with high slope, or underestimates deforestation at locations with low slope).
Figure 11: Brazilian Amazon deforestation logistic regression coefficient 95% confidence intervals using $N=963548$ map product points and $n=1386$ ground truth points. The PPI methods result in the smallest confidence intervals, while stratification is not generally effective. For slope and distance from rivers, the map-only coefficient falls outside the confidence intervals for all methods that use ground truth points.
4.2 Tree cover linear regression
Problem Formulation
Lastly, we estimate linear regression coefficients for the association between 2021 tree cover percentage in the contiguous United States and several covariates: aridity index, elevation, and slope.
Given $n$ data points, the linear regression model is
$\mu=\theta_{\text{aridity}}\textit{x}_{\text{aridity}}+\theta_{\text{elevation}}\textit{x}_{\text{elevation}}+\theta_{\text{slope}}\textit{x}_{\text{slope}}$
where $\mu$ is a $n\times 1$ vector of predictions, $\textit{x}_{j}$ is a $n\times 1$ vector of covariate $j$, and $\theta_{j}$ is the scalar linear regression coefficient for covariate $j$. We wish to estimate the $\theta_{j}$. A one-unit increase in covariate $j$ corresponds to a $\theta_{j}$ change in the outcome variable.
We use $N=983238$ map product points sampled uniformly at random from the 2021 USFS Tree Canopy Cover (TCC) product [forest2023usfs] at 30 m resolution. The TCC product is derived from Landsat and Sentinel-2 top-of-atmosphere reflectance data, USDA NASS CDL data, and terrain data. Its tree cover percentage predictions are obtained from random forest models trained on reference data from human-labeled high-resolution imagery. For ground truth, we manually label $n=983$ high-resolution satellite images from Google Maps for tree cover percentage, at locations sampled uniformly at random from the study area. Because the source of ground truth is from a uniform random sample, we do not use any stratified methods.
Figure 12: US tree cover linear regression datasets. We regress the tree cover percentage target variable (left) against three covariates (right): aridity index, elevation, and slope. We use the USFS Tree Canopy Cover map product and $n=983$ uniformly randomly sampled ground truth points.
Results
Figure 13: US tree cover linear regression coefficient 95% confidence intervals using $N=983238$ map product points and $n=983$ ground truth points. PPI outperforms the classical method. For elevation, the map-only coefficient falls outside the PPI confidence interval.
The 95% confidence intervals for the four linear regression coefficients are shown in Figure 13.
The PPI confidence intervals for coefficients are as follows. Tree cover percentage increases by 35 to 40 percentage points for each one-unit increase in aridity index, decreases by 2.4 to 4.8 percentage points for each one-kilometer increase in elevation, and increases by 0.9 to 1.3 percentage points for each one-degree increase in slope. The PPI effective sample size outperforms the classical method by a factor of 3.3 for aridity index, 2.4 for elevation, and 3.6 for slope. As in the other use cases, PPI uses the map product along with the ground truth to reduce uncertainty in the coefficient estimates compared to only using ground truth.
For elevation, the map-only coefficient computed from the map product ($-5.3$) falls outside the PPI confidence interval. This implies that there is bias in the tree cover map product’s association with elevation (e.g., the map product overestimates tree cover at low elevation locations or underestimates tree cover at high elevation locations). Similar to the deforestation logistic regression use case, using only the map product could result in misleading conclusions. If we had used only the remotely sensed tree cover predictions, we would have overestimated the magnitude of the relationship between elevation and tree cover due to map errors.
5 Discussion
Using remote sensing-based maps to draw scientific inferences requires uncertainty quantification and propagation of uncertainty to downstream analyses — otherwise, we risk making, in the words of [mcroberts2011satellite], just pretty pictures. In this work, we applied several methods that use remote sensing map products, along with a small amount of ground truth data, to estimate area and regression coefficients while correctly taking into account the errors in the map products. We find that stratified, post-stratified, and prediction-powered inference estimates result in lower uncertainty than classical estimates using only ground truth data sampled. Unlike map-only estimates using the map product alone, these estimates are guaranteed to be unbiased.
@@ 5 Discussion
Our experiments illustrated that a map-only estimate of area or regression coefficients from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias. In our simulation estimating maize fraction, map-only estimates deviated from the true maize fraction of 0.35 as the bias of the crop type map increased. Furthermore, accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
In our experiments on real data, the map-only estimate for deforestation in the Brazilian Amazon fell within stratified and PPI confidence intervals, suggesting an overall low-bias map product. However, the map-only estimates of the logistic regression coefficients on distance from rivers and slope fell outside the intervals computed through the classical, stratified, PPI, and stratified PPI methods. The map-only estimator would have concluded that distance from rivers is significantly more predictive of deforestation than it actually is, and that slope has no relationship to deforestation when it actually does. Thus, even when a map product is unbiased for the average outcome, it can be biased for the purpose of estimating regression coefficients. Similarly, the map-only estimate of the elevation coefficient in our linear regression for tree cover also overestimates the effect of elevation compared to PPI.
Effect of map accuracy.
How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the classical estimator depends on the quality of the map product. In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum. In PPI (and the post-stratified method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low. In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by 1.2×1.2$\times$1.2 ×.
@@
Comparing stratified, post-stratified, and PPI estimators, which estimator yields the smallest confidence intervals depends on how the strata were designed and the quality of the map product. Recall Table 2, which summarizes the conditions under which each method is most appropriate. Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI. This is a result of well-designed strata based on a remote sensing map of forest disturbance categories. PPI and the post-stratified estimator yielded comparable interval widths to each other. This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix A).
The picture changes, however, when we reuse the same stratified samples to estimate regression coefficients in regressions that the strata were not designed for. When we associated deforestation with covariates in a logistic regression, PPI and stratified PPI resulted in narrower confidence intervals than the stratified estimator across all 4 coefficients (distance from road, distance from river, elevation, and slope). The worse performance of the stratified estimator is because the strata were designed to minimize the variance of deforestation within each strata, which does not necessarily correspond to better regression estimation. Indeed, the stratified estimator was actually worse than the classical estimator for certain coefficients.
@@
An advantage of PPI compared to the stratified estimator, then, is its flexibility. While the stratified estimator requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset. Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain confidence intervals may result in the best performance across a variety of downstream use cases. Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
Recommendations for map producers and users.
Ultimately, the best methods combine map products with ground truth to yield the narrowest intervals (highest certainty). We recommend that producers of machine learning-generated remote sensing maps use a (simple or stratified) randomly sampled holdout ground truth dataset for uncertainty quantification. The criterion for this holdout set is that the model, including during hyperparameter tuning, was not trained on it; a standard test set would be adequate, provided that it is randomly sampled. Map producers should release this ground truth data alongside their map product to make their map suitable for downstream scientific inferences. In the absence of a provided holdout set (currently, releasing ground truth data is not the norm), we recommend that map users generate their own ground truth data for the variable of interest if possible.
Limitations.
All of the estimation methods described in this paper (except for map-only estimators) are limited to areas in which ground truth data is available. However, this would be true for any assumption-free, data-driven method that quantifies uncertainty and corrects for map bias. Furthermore, we can only ensure that our estimates are unbiased when the ground truth points are randomly sampled. Together, these attributes mean that such methods may be especially straightforward to use when ground truth data can be generated de novo, such as from human inspection of available satellite imagery, and more challenging to use when ground truth data requires bespoke surveys. Another limitation is that our estimates only quantify uncertainty in the target variable map product, and do not account for noise in either the ground truth data or (in the regression use cases) the covariates. We leave these investigations to future work.
6 Acknowledgments
[…]
Appendix A Mathematical comparison of PPI and post-stratified area estimator

### v4

@@ E.3 Area estimation examples
We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our area estimation use cases. The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points. The confidence intervals are unbiased and guaranteed to have the appropriate (e.g., 95%) coverage of the true area fraction even with this data-driven tuning [angelopoulos2023ppi++].
Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
E.3.1 Simulation: Maize area in Iowa
@@
In each trial, we sample $n=100$ “ground truth” points uniformly at random from the $N=100000$ map product points. Initially, we set the map product and ground truth equal to each other at the 100 ground truth locations. We then experiment with degrading the map product quality as follows. We systematically add noise or bias to the map product points while keeping the ground truth points constant, and then compare our maize fraction estimates across the map-only (pixel-counting), GT-only, and PPI estimators. We average results over 100 trials.
@@
We define adding “unbiased noise” at level $p$ to the map product as resampling each of the 100000 map product points with probability $p$ from the Bernoulli(0.35) distribution. In other words, about $p$ fraction of the points are randomly labeled, while the remaining $1-p$ points are unchanged. Each resampled point is labeled as maize with probability 0.35 and non-maize with probability 0.65. We experiment with noise probabilities $p$ from 0 to 1 at increments of 0.1.
@@ Results
As we add increasing amounts of random noise to the map, all confidence intervals for maize area remain centered at the true maize fraction of 0.35 (Figure 9). This is what we expect, since our noise-adding process resamples from the Bernoulli(0.35) distribution, causing the overall fraction of maize to remain the same. The GT-only confidence interval stays constant because it only uses ground truth points. The map-only confidence intervals are small because $N=100000$ is large, but they do not account for error in the map product. However, the map-only confidence intervals remain unbiased because the map still has true maize fraction close to 0.35 regardless of noise level.
When the noise probability is low, the PPI confidence intervals are significantly smaller than the GT-only intervals. This is because the error of the map product ($\hat{Y}-Y$) is small at the ground truth locations, so the variance of this error is also small, resulting in a small PPI standard error. As noise increases, the variance of the error of the map product increases; the PPI confidence intervals grow wider, eventually converging to the GT-only confidence interval at noise probability $p=1$. (At $p=1$, the map is complete noise and therefore does not provide any information for estimating the area beyond the information contained in the ground truth sample.) Consequently, the PPI effective sample size decreases from $n_{\text{effective}}=578$ when $p=0.1$ to $n_{\text{effective}}=100$ when $p=1$ (which matches the number of actual ground truth points $n=100$). Thus, the effectiveness of PPI is dependent on the quality of the map product. The post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI.
@@
The estimated maize fraction confidence intervals after adding bias to the map product are shown in Figure 10. Again, the GT-only confidence interval stays constant and remains centered at the true maize fraction because it only uses the ground truth points. As bias increases, the fraction of map product points labeled as “maize” increases. As a result, the map-only estimates increase with the bias value, and generally do not remain centered at the true maize fraction of 0.35. The map-only confidence intervals are small because $N=100000$ is large, but they are unreliable because they do not correct for the bias in the map product. In contrast, the PPI estimates remain centered at the true maize fraction because PPI corrects for bias in the map product. The PPI confidence intervals are smaller than the GT-only intervals because the variance of the map product error $\hat{Y}-Y$ is small (since only about 40% of the map product points are resampled, regardless of bias level). Thus, PPI provides unbiased estimates with lower uncertainty than using only the ground truth points. Again, the post-stratified confidence intervals and effective sample sizes are nearly identical to those of PPI, as expected.
We also plot the maize fraction confidence intervals against overall map accuracy for both the noise and bias experiments (Figure 11). As noise increases, map accuracy decreases, so low map accuracy is associated with wide PPI and post-stratified confidence intervals. Thus, the left panel of Figure 11 resembles an inverted version of the left panel of Figure 9.
Figure 11: Simulation estimating maize fraction 95% confidence intervals at different map product overall accuracy levels obtained by varying noise (left) and bias (right), using $N=100000$ map product points and $n=100$ “ground truth” points. Post-stratified and PPI confidence intervals are overlapping.
As bias increases and the fraction of map product points labeled as “maize” increases, map accuracy decreases linearly. (To see this, note that if each resampled point is labeled as maize with probability $p$, then an expected $0.35(p)+0.65(1-p)=0.65-0.3p$ fraction of the resampled points will match the ground truth.) Thus, the map-only estimates of maize fraction decrease linearly as map accuracy increases. In the right panel of Figure 11, a map accuracy of 0.86 corresponds to a map-only maize fraction estimate of 0.21, significantly lower than the true fraction of 0.35. This shows that even a map with high map accuracy can yield biased estimates under the map-only (pixel-counting) estimator.
E.3.2 Deforestation area in Brazilian Amazon
@@
We create a binary deforestation map product using all $N=4.4\times 10^{9}$ pixels from the NASA Global Forest Cover Change (GFCC) map [townshend2016global] (at 30 m resolution) in the study area (Figure 12). The GFCC product is derived from Landsat 5 and Landsat 7 surface reflectance imagery, and includes maps for tree cover percentage in each pixel in five-year increments from 2000 to 2015. We consider a map pixel deforested if its canopy cover percentage satisfies $\text{canopy}_{2015}-\text{canopy}_{2000}\leq-25\%$ (Figure 6 in Appendix C).
We use a ground truth Amazon deforestation and forest disturbance dataset created by [bullock2020satellite] using time-series analysis of Landsat imagery and high-resolution imagery from Google Earth between 1995 and 2017. The dataset records the years corresponding to deforestation, degradation, and natural disturbance events at each ground truth point. We use this to construct a binary ground truth dataset restricted to Brazil, in which we consider a point deforested if it experienced a deforestation event between 2000 and 2015. The original dataset is obtained by stratified random sampling, using a stratification map (also created by [bullock2020satellite]) that contains 7 strata related to deforestation status. To generate a GT-only estimate baseline, we construct a “uniformly random sampled” ground truth dataset by downsampling the points in each of the strata to match its area proportion. For a fair comparison, we also downsample the original stratified ground truth dataset to match the size of the “uniformly random” dataset. This results in two ground truth datasets of $n=1386$ points each, corresponding to uniform random sampling and stratified random sampling. Both ground truth datasets are used with PPI to obtain confidence intervals.
@@ Results
Figure 13: Brazilian Amazon deforestation 95% confidence intervals using $N=4.4\times 10^{9}$ map product points and $n=1386$ ground truth points. The stratified, post-stratified, and PPI estimators all result in smaller confidence intervals for estimated deforestation than GT-only estimation, with the stratified estimator and stratified PPI having the best performances.
Each method results in a point estimate that between 5% to 6% of the Brazilian Amazon was deforested between 2000 and 2015. The stratified, post-stratified, and PPI estimators all result in smaller 95% confidence intervals for estimated deforestation than GT-only estimation (Figure 13). The post-stratified and PPI confidence intervals are fairly similar, with respective effective sample sizes that are $1.9$ and $1.8$ times the actual number of ground truth sample points. The stratified and stratified PPI confidence intervals have the best performance, with effective sample sizes that outperform the GT-only method by a factor of about $2.4$. (Put differently, stratified PPI has effective sample size $3382$, so using the remotely sensed deforestation map added the equivalent of $3382-1386=1996$ ground truth sample points.) Thus, the methods that use both map product and ground truth reduce uncertainty in our estimates, compared to using only ground truth. In particular, the stratified estimator is effective because the strata map is designed so that points within each stratum have low variance in deforestation status.
@@
E.4 Discussion
@@
Our experiments illustrated that a map-only area estimate from a map without bias correction can yield incorrect inferences, and we differentiated between the influence of noise versus bias. In our simulation estimating maize fraction, map-only estimates deviated from the true maize fraction of 0.35 as the bias of the crop type map increased. Furthermore, map accuracy is not necessarily a good proxy for bias; even at seemingly high map accuracies ($\text{accuracy}>0.85$), the map-only approach estimated maize fraction at 0.21 (a 40% underestimate).
@@
How much stratified, post-stratified, and PPI estimators reduce uncertainty relative to the GT-only estimator depends on the quality of the map product. In stratification, the strata map should be designed such that the variance of the quantity of interest is low within each stratum. In PPI (and the post-stratified method), the variance of the difference between the map product predictions $\hat{Y}$ and the ground truth values $Y$ should be low. In our maize fraction experiments, a highly accurate map ($\text{accuracy}=0.96$) was equivalent to having nearly 6×6$\times$6 × as much ground truth data, whereas a low-accuracy map ($\text{accuracy}=0.73$) only increased the effective sample size by 1.2×1.2$\times$1.2 ×.
@@
Empirically, when estimating deforestation fraction in the Brazilian Amazon, we observed the smallest confidence intervals when stratification was used alone or in conjunction with PPI. This is a result of well-designed strata based on a remote sensing map of forest disturbance categories. PPI and the post-stratified estimator yielded comparable interval widths to each other. This is expected, as the post-stratified estimator yields similar confidence intervals as PPI when $N$ is large (Appendix F).
@@
An advantage of PPI compared to the stratified estimator is its flexibility. While the stratified estimator requires us to choose a map product before sampling, PPI can be applied to any map product and any randomly sampled ground truth dataset. Given the expensive nature of ground truth data collection, sampling without choosing a strata map and then using PPI to obtain confidence intervals may result in the best performance across a variety of downstream use cases. Or, if stratified samples are already available from area estimation, they can be combined with PPI to estimate regression coefficients.
Appendix F Mathematical comparison of PPI and post-stratified area estimator